# Generalized Random Forests using Fixed-Point Trees

### v3

@@ Appendix E Additional Examples of Pseudo-Outcomes
In this special case, we recover the conventional splitting algorithm used for univariate responses (Breiman et al., 1984; Breiman, 2001) or for multivariate responses (De’ath, 2002; Segal, 1992). Trees grown using $\rho_{i}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}$, $\rho_{i}^{\texttt{grad}}$, or $Y_{i}$ will identical to one another because CART splits are scale and translation invariant with respect to the response.
@@ Appendix F Real Data Comparison: California Housing
where we suppose that our effects are local to spatial coordinates $x:=(\texttt{latitude}_{i},\texttt{longitude}_{i})$, $Y_{i}$ denotes the log median housing value of the census block, and the primary regressors $W_{i}=(W_{i,1},\ldots,W_{i,6})$ passed to the model were
  • median housing age,
  • log(total rooms),
  • log(total bedrooms),
  • log(population),
  • log(households),
  • log(median income),
and where $\theta^{*}_{k}(x)$ denotes the geographically-varying effect of the corresponding regressor $W_{i,k}$, for $k=1,\ldots,6$. The empirical distribution of the transformed regressors passed to each of the GRF models is seen in Figure 21.
@@
We target GRF estimates $\hat{\theta}(x)=(\hat{\theta}_{1}(x),\ldots,\hat{\theta}_{6}(x))^{\top}$ of $\theta^{*}(x)=(\theta^{*}_{1}(x),\ldots,\theta^{*}_{6}(x))^{\top}$ based on the GRF-FPT algorithm with approximate OLS coefficients described in Section 5, and compare GRF-FPT estimates with those given by GRF-grad. Forests were fit using the grf::lm_forest function, which trains a forest and solves for the GRF estimates $\hat{\theta}(x)$ based on a varying coefficient model (37). All versions fit a forest of 2000 trees, the default settings of the original R implementation (Tibshirani et al., 2024), and a target minimum node size of 5 observations.
Results.
Table 4 summarizes the computational benefit of GRF-FPT applied to the California housing data. Figures 5, and 20 illustrate the local estimates $\hat{\theta}(x)$ made by GRF under the FPT and grad algorithms, respectively.
@@ Results.
Figure 20: Geographically-varying local estimates $\hat{\theta}(x)=(\hat{\theta}_{1}(x),\ldots,\hat{\theta}_{6}(x))$, solved under the GRF-grad algorithm.
Figure 21: Empirical distribution of the regressors passed to each GRF algorithm.

### v4

@@ E.1 Pseudo-outcomes for nonparametric regression
In this special case, we recover the conventional splitting algorithm used for univariate responses (Breiman et al., 1984; Breiman, 2001) or for multivariate responses (De’ath, 2002; Segal, 1992). Trees grown using $\rho_{i}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}$, $\rho_{i}^{\texttt{grad}}$, or $Y_{i}$ will be identical to one another because CART splits are scale and translation invariant with respect to the response.
@@ Appendix F Real Data Comparison: California Housing
where we suppose that our effects are local to spatial coordinates $x:=(\texttt{latitude}_{i},\texttt{longitude}_{i})$, $Y_{i}$ denotes the log median housing value of the census block, and the primary regressors $W_{i}=(W_{i,1},\ldots,W_{i,6})$ passed to the model were median housing age, log(total rooms), log(total bedrooms), log(population), log(households), and log(median income). Here, each $\theta^{*}_{k}(x)$ denotes the geographically-varying effect of the corresponding regressor $W_{i,k}$, for $k=1,\ldots,6$. The empirical distribution of the transformed regressors passed to each of the GRF models is seen in Figure 17.
@@
We target GRF estimates $\hat{\theta}(x)=(\hat{\theta}_{1}(x),\ldots,\hat{\theta}_{6}(x))^{\top}$ of $\theta^{*}(x)=(\theta^{*}_{1}(x),\ldots,\theta^{*}_{6}(x))^{\top}$ based on the GRF-FPT1 and GRF-FPT2 algorithms described in Section 6, and compare those to GRF-grad. All forests were fit using the grf::lm_forest function, which trains the Stage I forest and optionally solves for the Stage II estimates $\hat{\theta}(x)$ for varying coefficient models (65). All versions fit a forest of 2000 trees, the default settings of the original R implementation (Tibshirani et al., 2024), a subsample ratio of $0.5$, and a target minimum node size of 5 observations.
@@
Results.
Table 7 summarizes the computational benefit of GRF-FPT applied to the California housing data. Figures 4 illustrates the local estimates $\hat{\theta}(x)$ made by GRF-FPT2, while Figure 16 illustrates the fits under GRF-FPT1 and GRF-grad.
Refer to caption Refer to caption
Figure 16: Geographically-varying local estimates $\hat{\theta}(x)=(\hat{\theta}_{1}(x),\ldots,\hat{\theta}_{6}(x))$, fit under GRF-FPT1 (top) and GRF-grad (bottom). Results for GRF-FPT2 are presented in Figure 4 found in Section 7.
Figure 17: Empirical distribution of the regressors from the California housing data passed to GRF.